# A Perspective on Foundation Models for the Electric Power Grid

### v1

@@ 3.3 Hurdles to implementation
4 A concept of foundation model for the power grid
In the previous section, we outlined the salient qualities of FMs by pointing to their established applications in language and weather. Here, we discuss how the challenges in the future power system open up new opportunities for FMs. FMs can reduce computational burdens, accelerating simulations and thus enabling the transition to a sustainable and reliable power system. These models excel in handling complex power grid tasks in near real-time and provide more accurate forecasting or reconstruction of key variables than classical methods. Such problems are well-suited for FMs as they fundamentally involve token prediction or reconstruction tasks. Additionally, an ML agent can be trained to act as an expert in decision-making processes, such as grid expansion, network reconfiguration, and other tasks that often require computationally intensive solutions. Here, we explore the potential implications of developing a GridFM, examining the important components and steps to build such a model.
4.1 Data preprocessing and token design
The successful construction of FMs requires the careful design of tokens to represent data elements effectively. Word and image tokens are essential for successful natural language processing and computer vision training. While these domains have established tokenization methods, strategies must be explored for other domains. For example, graph-based models define a graph vocabulary [67] to capture structural and semantic properties, while time series models segment data into subseries-level patches [68].
Designing tokenization strategies for non-text and non-visual power grid data can be difficult, because the grid involves very heterogeneous, multi-modal datasets, sensor time series, categorical equipment status data, and spatial grid topology data. The PyPSA-Eur dataset [69] is an excellent example of this complexity, where researchers provide a platform to access multi-modal data for the European power system, i.e., network data, electrical demand time series, renewable time series, and geographical potential for solar and wind generations. A promising tokenization strategy for multi-modal data was demonstrated for vision tasks by the massively multi-modal masked modeling architecture (4M) [70]. In this work, good scalability was achieved by unifying the representation space of all modalities, by mapping them into discrete tokens, and performing multi-modal masked modeling on a small randomized subset of tokens. In addition to tokenization, other preprocessing techniques are essential for training models on grid data. For example, normalization methods are critical for training time series models to avoid distribution shift issues [68, 71]. Nevertheless, the alarm data in power grids resembles tokens and natural language elements [72], making FMs applicable for forecasting alarm sequences and filling in missing alarm data.
4.2 Model architecture
Our GridFM model must work with data from at least four modalities: tabular, time series, graph (static and spatiotemporal), and geospatial. Additionally, we expect the power grid dynamics to require a similar number of trainable parameters than weather and climate systems, which are also governed by highly non-linear and complex physical laws. Therefore, an estimate of the number of trainable parameters for GridFMs can be inferred from the FMs developed for these systems, which currently have a few hundred million parameters [53, 73].
We envision two designs for a GridFM that account for the complexity of the grid and its varying transmission and distribution network properties, covering system dynamics from operations to planning. Firstly, it can be designed as a single multi-modal spatiotemporal model. Alternatively, it can consist of multiple FMs integrated into a mixture-of-experts (MoE) framework [74].
The topology of the power grid and its physical properties align well with the message-passing mechanisms used in graph neural networks (GNNs). GNNs are designed for graph-structured data and, through their message-passing mechanisms, learn representations of graph structures called embeddings. These embeddings are learned by iteratively aggregating the features of neighboring nodes. Unlike other deep learning methods that either fail to generalize or require complex data preprocessing, GNNs can generalize to different grid network topologies. Selecting a suitable GNN architecture, however, is challenging. It was argued that the power grid’s unique features necessitate GNNs with depths of up to 27 layers to effectively solve the power flow problem  [75]. More recently, on the other hand, it was demonstrated that more shallow GNNs can achieve very good results for power flow, optimal power flow, and cascading failure analysis on various benchmark transmission grids [13]. Specifically, it was shown that graph attention networks  [76] and transformer-based models perform the best. Attention-based models on graphs assign the importance of a node to its neighbors by learning a score for each node pair, which allows for a more effective understanding of temporal and geographical patterns. Furthermore, attention-based methods such as GraphAny [77] demonstrate high transferability and generalizability to unseen graph topologies for foundation models. Given these insights into the effectiveness of GNNs for power grid analysis, exploring more advanced modeling strategies becomes essential to capture the complexities of the multi-level structure of the power grid. A promising strategy is a hierarchical model architecture, e.g., using GNNs and sublevel time series modeling [68]. Hierarchical GNNs [78] allow for modeling transmission and distribution-level grid topologies. Sublevel time series models can train the FM with various temporal dynamics and resolutions. In addition, hierarchical GNNs can be combined with geometric deep learning for spatiotemporal graphs [79]. The combined use of these sophisticated approaches requires further exploration. It is important to note that the graph representations can go beyond the physical grid topology [80].
Here, we propose a concept of a multi-modal FM for power grids as shown in Figure 4. The model reads structured graphs in which the grid topology is mapped to nodes and edges using GNNs as encoders. Inputs serve as embeddings, including static (tabular and static graphs) and dynamic data (time series, spatiotemporal graphs, and geospatial data). The model also reads the input from variations of the same or different power grids. The input data is partially masked to train an auto-encoder for reconstructing the inputs. The model outputs the reconstructed inputs: measured signals, topology, and simulated data. While using a self-supervised learning approach to reconstruct signals is established, topology reconstruction entails novel techniques. A promising candidate for such a task is the integration of a link prediction approach into the GridFM model. Such an approach allows the model to learn the presence of a masked link using the observed (masked) graph topology [81].
Figure 4: The GridFM concept, which reconstructs power grid-relevant data. Left: Input. Middle: Model Architecture. Right: Output. Grayed-out regions indicate masked tokens. We present examples of inputs to be reconstructed by the foundation model, including time-series signals at the bus, branch, and system levels, tabular data from load flow simulations, and topology reconstructions. Such a model aims to handle various configurations of the same power grid and entirely different power grids.
4.3 Downstream tasks
Once trained, the GridFM can address multiple downstream tasks, with or without additional encoding layers or fine-tuning. We summarize in Figure 5 and discuss the following eight classes of such tasks.
Figure 5: Downstream tasks: Examples of power system problems that a GridFM can address.
Transient and dynamic stability
Here, the GridFM is employed to estimate how a system may respond to disturbances, analyzing signals in both time and frequency domain. The GridFM, when fine-tuned to reconstruct such signals, can help with missing data or predictive modeling. Foundation models with physics-informed learning can efficiently solve the differential algebraic equations governing rotor angle and voltage stability [82]. This can significantly speed up the computation of stability margins and help identify potential stability issues. Moreover, the rapid computation of fault currents using Foundation Models can aid operators in accurately localizing and classifying faults, enabling faster system restoration and minimizing the impact of disturbances on the power system [83].
Load flow
In power system operation and planning, load flow analysis determines the electrical state of the power system, e.g., voltage and phase angles at buses and power flows in branches. These analyses benefit from AI/ML methods, significantly decreasing computational burdens [13]. Traditionally, the power flow and optimal power flow problems are solved with iterative techniques. Due to their non-linear nature and difficulties in converging for large problems, simplifications are often used. Scientific ML techniques have been shown to provide solutions without employing simplified equations, competing with traditional methods in terms of accuracy while being drastically faster [84]. Therefore, with a significant computational speed-up, GridFM has great potential to provide more accurate and scalable solutions for power systems [82].
@@ 4.3 Downstream tasks
Forecasting is used in power grids for multiple tasks, including demand prediction, renewable production forecasting, and electricity price prediction. Demand and renewable forecasting are vital in ensuring power generation-demand balance, which is becoming increasingly important in a grid with high renewable penetration. The self-supervised learning paradigm used in FMs and their predictive ability make them intrinsically suitable for forecasting [85]. Moreover, FMs can integrate exogenous variables, such as numerical weather predictions, which also carry associated forecast uncertainties.
@@
Here, the concerns include frequency and voltage control, network reconfiguration management, protection activation, and load shedding. FMs will help optimize and control the power system through sequential decision-making agents [86]. A GridFM can enhance reinforcement learning (RL) agents by introducing prior knowledge of the environment before and during training, improving learning efficiency, and reducing risks.
Electricity markets
Electricity markets will benefit from the computational efficiency and predictive ability of a GridFM. The different markets, including futures, balancing, day-ahead, and intraday markets, can be optimized using the base GridFM or fine-tuned versions. Furthermore, the model will identify the security-based re-dispatch after the day-ahead market clearing in a timely manner [87]. It will also help to pinpoint the location, amount, and type of ancillary services from distributed sources for congestion management in the future power grid with high shares of renewables. Finally, a GridFM will provide prompt solutions for economic dispatch and unit commitment problems, traditionally solved as linear and mixed-integer optimization problems [88].
System security
ML methods have been used to identify faulty components [12] and to speed up simulations like $N-1$, $N-K$ contingency, and cascading failure analyses [89], all critical tasks for operation and planning. Such models are faster and better leverage past knowledge for security predictions than traditional tools. Nevertheless, these analyses require additional encoding to incorporate topology with bus or branch-level features to inform about the final state of the entire grid. In this context, a GridFM tuned for system security will explore a broader and more complete set of scenarios, and thus both aid in designing a reliable power system and provide a system security tool for power system operations planning.
@@
Rapid electrification of the industrial and automotive sectors and climate change’s impact on generation profiles make expansion planning a crucial task at distribution and transmission levels. The generative capabilities of a GridFM to create future scenarios for load and generation (e.g., using public projections of electric vehicle sales and national renewable energy plans) will enhance decision-making. By integrating scenario generation with graph generation and deep reinforcement learning techniques, optimized and upgraded power systems can be designed to capture a broader spectrum of future extreme scenarios [90]. Additionally, by exploiting cross-country information (e.g., the impact of interconnection flows), a GridFM will improve expansion planning strategies to prepare for these effects and enable the energy transition in the following decades.
@@
Cybersecurity challenges could be tackled by leveraging FMs to speed up power system simulations. These FMs facilitate risk assessment and incident response in transmission and distribution systems by identifying and prioritizing threat scenarios [91] and designing optimal responses to maintain grid reliability [92]. They can generate synthetic cybersecurity data by learning complex power system dynamics to overcome data scarcity, improving intrusion detection, and enabling dynamic stress testing against intelligent adversaries [93]. A cyber-physical FM can combine SCADA and AMI data with cybersecurity information to enhance cyber threat mitigation, detection, and response.
5 Conclusion
@@
The rapid emergence of FMs is reshaping the AI landscape. While such models are well established for text and image analysis, here we have introduced the idea of a GridFM: an FM for the power grid, able to learn from time-series, geospatial, and other data to improve power grid planning, operations, and control. We first motivated the need for a GridFM by pointing to the growing gaps between computational capabilities and the industry’s need to cope with the increasing uncertainty and complexity of the power grid, largely due to the energy transition. We then reviewed the three core advantages of the FM approach for power systems: excellent missing token prediction skills (signal reconstruction); broad adaptability and scalability; and simulation acceleration. In another domain characterized by apparently intractable complexity, namely weather forecasting, we have seen AI weather FMs approximate the fundamental equations governing weather to produce much more computationally efficient models (more than 4 orders of magnitude speed-up).
In a similar manner, we propose to develop a GridFM that learns the behaviors and dynamics from measurements and simulations for different network configurations. The superior computational performance of such a GridFM has the potential to accelerate grid analyses by several orders of magnitudes. This acceleration will, in turn, enable modeling of the increasingly complex (e.g., bidirectional power flow, combined transmission, and distribution) and uncertain grid (e.g., by modeling different scenarios going beyond the traditional N-1 analysis). The promise of such a GridFM is that once pre-trained on multiple grids and various networks, it can be readily fine-tuned to specific grid topologies (distribution and transmission). In summary, we argue that research in the proposed FM for the power grid is imperative to facilitate the energy transition. We envision the power system and AI communities collaborating to take on this “moonshot” and thus revolutionize power grid analysis, leveraging AI advancements to meet common goals and needs.
Acknowledgement
The work was supported by the respective organizations of the authors. The authors acknowledge the invaluable discussions and interactions in the working group for Foundation Models for the Power Grid, including Roya Ahmadi, Rich Argentieri, Gianni Vittorio Armani, Venkat Balachandran, Ranjini Bangalore, Venkat Banunarayanan, Edmon Begoli, Larry Bekkedahl, Keith Benes, Michelle Berti, Budhu Bhaduri, Roberto Bianchessi, Gilbert Bindewald, Louise Blais, David Bromberg, Matthias Bucher, Eren Cam, Supriya Chinthavali, Seong Choi, Julien Chosson, Joseph Collum, Magnus Cormack, Peter De Bock, Fei Ding, Kathy Duviella, Darlain Edeme, Yamshid Farhat, Mark Fenton, Marcus Freitag, Alirez Ghassemian, Tassos Golnas, Gavin Goodland, Luca Grella, John Grosh, Benois Grossin, Andy Hackett, Hagen Haentsch, Zeryai Hagos, Grahame Hay, Frank Hellmann, Andy Henton, Raffael Hilber, Dan Hilton, Gabriela Hug, Alex Hurley, Sorana Ionescu, Kevin Johnston, Gareth Jones, Adrian Kelly, Bernd Klockl, Teddy Kott, Benjamin Kroposki, Teja Kuruganti, Larry Light, Dalton Lunga, Matt Magill, Adrian Maldonado, Antoine Maret, Pete Marsh, Russel Mason, James McClean, John Messer, Nikki Militello, Jordan Murkin, Xabi Muruaga, Indu Nambiar, Christian Nauck, Caleb Northrop, Gabby Nyirjesy , Marco Padula, Patrick Panciatici, Raul Alfaro Pelico, Mark Petri, Pierre Pinson, Abhishek Pontis, Josh Power, Alan Pressman, Feng Qiu, Mark Rafferty, Pippa Robertson, Nicoletta Rocca, Caroline Rose-Newport, Lyndon Ruff, Bryan Sacks, Kelly Sanders, Giovanni Sansavini, Michael Sheppard, Robert Shorten, Vikas Singhvi, Stanislav Sobolevsky, Pratik Sonthalia, Brian Spears, Declan Stock, Olle Sundstroem, Matthew Tarduogno, Roger Thompson, Anna Carolina Tortora, Alex Vivas, Lucy Vu, Ben Weiner, Casey Werth, Daniel Wetzel, Gary White, Xin Wu, Steven Wysmuller, Yunchi Yang, Thomas Zadlo, Marzia Zafar and many others. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.
@@ Acknowledgement
Conceptualization, H.F.H., L.S.A.M. and J.W.; Investigation, H.F.H., T.B., B.G., L.S.A.M., A.P., A.V., J.W., J.B., A.B.M., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., and A.Z.; Writing – Original Draft, H.F.H., T.B., B.G., L.S.A.M., A.P., A.V., J.W., J.B., A.B.M., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., and A.Z.; Writing – Review & Editing, A.J.B., R.J.B., B.P.B., J.S. and S.S.; Project Administration, H.F.H.

### v2

@@ 3.3 Hurdles to implementation
4 Proposed implementations for GridFM
Figure 4: Applications: Examples of power system problems to be solved with GridFMs.
The broader vision of a GridFM is the enablement of all or as many applications as possible. However, even if initially GridFMs would only be suitable for a single application, the superior adaptability of FMs over bespoke AI models through fine-tuning to thousands of different utilities with their specific data and requirements justifies the FM approach.
In the above context, fine-tuning can mean different things. A GridFM may be fine-tuned to a set of diverse applications like load forecasting, fault detection and others. Initially, this may be too ambitious and challenging, considering the diversity of downstream applications as shown in Figure 4, each requiring different data, temporal scales, and pre-training strategies. But fine-tuning may also refer to tailoring a GridFM for a given application with user-specific (and potentially private) data to the particular user needs and requirements. For example, just in the United States, there are 3,200 utilities [71], with their own private and siloed data and grid-specific requirements. These utilities may not have sufficient data by themselves but could directly profit from the proposed approach. This multi-level fine-tuning approach is not unlike what can be observed in geospatial or weather foundation models, where FMs have shown excellent adaptability to different geographic regions, spatial and temporal resolutions for a given downstream application [72, 73, 50].
We thus propose an implementation road map as outlined in Figure 5, with a first, more specific version of GridFMs, pre-trained for power flow applications (GridFM–v0), which is expected to provide real-world benefits in the next 1-3 year. In the long-term, we envision expanding and addressing more grid applications by including more diverse data modalities to pre-train a single GridFM or by combining different GridFMs, as discussed in Section 4.2. The rationale to start with power flow is as follows:
(1) A highly relevant pre-training task
FMs are known to excel when the pre-training task is closely related to the enabled downstream applications [74]. Consequently, with AC power flow at the core of power system analysis [75], its reconstruction is an ideal pre-training task for GridFM–v0. This enables power flow related applications in the area of load flow (optimal power flow, state estimation, power flow), system security (contingency analysis, cascading failure), and beyond.
(2) AI power flow models are computationally more efficient
Current power flow solvers, e.g. using Newton-Raphson methods, are computationally expensive, especially for large grids. For instance, for an N-1 contingency analysis on the Western Interconnection, it takes approximately 5 minutes per scenario [76], while meeting the NERC TOP-001-3 requires running these analyses within a limited time period. In contrast, GridFM–v0 is projected to achieve a computational speed-up of 3 to 4 orders of magnitude over conventional solvers, as demonstrated on IEEE grid cases with up to 118 buses [77]. For larger grids, we anticipate a similar or greater speedup, as GNNs exhibit linear complexity in the number of grid buses [78], while Newton-Raphson methods scale quadratically with the number of buses [79]. This significant efficiency gain will enable real-time analysis of multiple scenarios of more complex grids [77], which will become crucial for maintaining system stability and optimizing future grids under uncertainties stemming from the energy transition and from climate change [76].
(3) There have been significant advances in GNN-based power flow
Using GNNs to represent the power grid and mapping power flow to the message-passing scheme of GNNs has been successfully demonstrated recently [77]. These models replicated the results of traditional power flow solver with less than 1% error on the predicted quantities, despite past GNN training challenges [80]. While these models were trained in a supervised way, recent results of the previously discussed GNN-based weather FMs [72, 73, 50] indicate that learning power flow with GNNs in a self-supervised way is a valid approach for a GridFM. Furthermore, large-scale data foundation models (FM) based on GNNs have already been developed [48], and a comprehensive review can be found in [81].
Figure 5: GridFM implementation road map with four near-term phases to develop GridFM–v0, a GridFM for power flow related applications. The icons refer to academia, power system industry, and AI community. The long-term goal is creating a family of GridFMs for a broader range of grid challenges.
4.1 Near-term: GridFM–v0, a GridFM for power flow applications
The four phases as foreseen for the development of GridFM–v0, and as depicted in Figure 5 are discussed in detail in the following.
4.1.1 Phase 1: Identify, collect and generate training data
Training GridFM–v0 requires a large collection of solved power flow problems for different grid topologies, parameters, and different load conditions. In recent years, large power grid data-sets have been made available, created by (1) collecting grid topology and measurement data, (2) generating varying synthetic load conditions per grid model and (3) using established solvers to compute a wide range of power flow solutions under different operational conditions.We aim at using the datas-et provided by [13], which contains power flow solutions under real load conditions on 4 different grid topologies. Additionally, to complement this data, diverse power flow problems using traditional power flow solvers [82, 83, 84] will be computed, with power grid topologies from PGLIB-OPF [85] and standard IEEE benchmarks and with topology and load perturbations. Consequently, our model will be pre-trained on data from realistic topologies with a broad range of operational conditions. This approach facilitates the model to generalize well across topologies, as demonstrated in [86]. For model fine-tuning, data-sets from [86] for system security analyses and [87] for Optimal Power Flow (OPF) will be used. The latter contains 300,000 solved optimal power flow problems for ten different grid topologies each, including solutions for N-1 contingency analysis, with grid sizes ranging from 14 to 14,000 buses.
4.1.2 Phase 2: Model architecture development
Figure 6: GridFM–v0 is pre-trained to reconstruct masked power flow data from input graphs.
In the second phase, suitable architectures will be evaluated by pre-training performance analysis, and by adjusting models, training strategies, and loss functions. Figure 6 illustrates the Masked Auto-Encoder (MAE) pre-training task of GridFM–v0. Intuitively, the grid is modeled as a graph, where nodes represent electrical buses with active power, reactive power, voltage magnitude, and voltage angle, $(p_{i},q_{i},v_{i},\delta_{i})$, as node variables. We opted for this graph representation of power systems, but others can also be considered [88]. Transmission lines and transformers are modeled as graph edges. An autoencoder [89, 90] is trained in a self-supervised manner to reconstruct masked node features from power flow solutions, each corresponding to a specific topology and load condition. This amounts to solving the power flow equations. During training, the loss function ensures meaningful reconstruction of node variables within the physical constraints. The total loss $\mathcal{L}=\mathcal{L}_{SCE}+\mathcal{L}_{powerflow}$ combines the Scaled Cosine Error (SCE) [89], which minimizes the distance between the original and reconstructed graphs, and $\mathcal{L}_{powerflow}$, which ensures that the power flow equations are satisfied. Recent studies show that an approach using physics-informed methods is the most effective for solving OPF. For example, [91] demonstrated that a hybrid loss function provides high-fidelity approximations for large-scale OPF, while [92] confirmed that this approach best solves AC OPF by integrating data-driven learning with physical laws. FMs, as a concept, are designed to incorporate diverse methods, including GNNs and PINNs, into a unified framework, enabling them to generalize effectively and adapt to various tasks, such as solving OPF problems.
4.1.3 Phase 3: Model scaling and validation on industry data
Due to regulatory constraints, it is unlikely or at least challenging to gain full access to sensitive data from operational power grids [93, 94, 95]. To ensure the model’s accuracy and relevance in practical applications, it is therefore imperative that a GridFM be validated on operational data and real grid topologies. To address these challenges, an iterative approach is adopted. Our GridFM–v0 will initially be pre-trained and tested on openly available data. Subsequently, it will be re-trained and validated in collaboration with electricity utility partners behind the firewall in a secure compute environment. The results of these validations will guide us to generate more suitable synthetic training data, update the architecture, loss functions etc. so that the improved GridFM yields satisfactory results in real-world conditions once fine-tuned. This approach ensures that the GridFM remains performant and aligned with industry needs without exposing nor pre-training the model on proprietary data.
4.1.4 Phase 4: Implement power flow-based use cases
GridFMs are anticipated to achieve a computational speed-up of 3 to 4 orders of magnitude over traditional power flow solvers, as stated previously, which will directly benefit various applications:
(1) Complementing Numerical Power Flow Solvers
For applications that require solving power flow frequently, and where speed is critical, linear approximations (such as DC power flow [96]) are typically used. For these applications, faster AI-based power flow solvers offer a trade-off between “traditional” DC power flow and AC power flow solutions. We expect minimal fine-tuning of these models, as the task is equivalent to the suggested pre-training reconstructing task.
(2) Contingency Analysis
Contingency analysis is essential for assessing grid resilience. On a grid with 1000 lines, simulating the loss of each individual line requires solving 1,000 distinct power flow problems. However, when considering scenarios that involve the simultaneous loss of any two components, the number of power flow analyses increases dramatically to $\binom{1000}{2}=499,500$. By utilizing a power flow solver that is 1000 times faster, operators will be able to perform contingency analyses across a much larger set of scenarios. For example, solving 499,500 scenarios with AI-based methods would take roughly the same time as solving only about 500 scenarios using conventional physics-based solvers. Therefore, a GridFM–v0 will greatly enhance grid analyses and contribute to operating and planning under uncertainty.
The power grid is continuously evolving, therefore, the GridFM–v0 must be able to account for different network topologies and different operating conditions of a same grid. This can be accomplished by training the model on grids with topological perturbations, as already demonstrated by [97], or by additionally training the same model on entirely different grid topologies, as done in [86].
(3) Optimal Power Flow (OPF)
The authors in [97] trained a model on synthetic data and demonstrated an OPF speedup of 2 to 4 orders of magnitude while maintaining errors of less than 1% in the objective value. Solving the OPF translates into fine-tuning the GridFM for reconstructing the generator power to minimize generation costs. Reconstructed node variables must adhere to bounds (e.g., generator capacity constraints), making this problem challenging. However, GridFM–v0 can be hierarchically integrated into optimization schemes where generator power is iteratively adjusted, solution feasibility is checked, and other variables are determined using the power flow solver.
4.2 Long-term Vision: GridFMs for a broader set of applications in power systems
In Section 4.1, we explored an initial version of GridFM for power flow applications, already addressing important challenges of the industry. The long-term objective, however, is to address as many applications from Figure 4 as possible, using a single GridFM or a family thereof, fully exploiting FM capabilities. Here we discuss how to build new industry-relevant applications on top of GridFM–v0.
4.2.1 Extending GridFM–v0 to a GridFM capable of handling new applications
@@
GridFM–v0 was pre-trained to reconstruct power flow data for a single time step. This can be readily extended to pre-training to reconstruct power data over multiple consecutive time steps, similar to Weather FMs [73]. This enhancement will enable the forecasting of future grid states. By incorporating weather and climate data into the node embeddings, we will further improve the model’s ability to predict generation and anticipate weather-induced grid disturbances.
Transient and Dynamic Stability Analysis
In addition to training GridFMs on static power flow data, like in the case of GridFM–v0, the pre-training can additionally involve reconstructing high-resolution real-time data obtained from PMUs, which enables effective transient state estimation. This approach will facilitate rapid fault current detection and localization, efficient system restoration, and disturbance mitigation [98]. Additionally, the generator rotor angle and voltage differential-algebraic equations can be added to the model using physics-based machine learning techniques [99] to, e.g., assess the grid’s stability margins.
@@
There are two concrete ways in which GridFMs will help in control operations. First, the generative capabilities of FMs as a conditional load and generation scenario generator, can be used to construct scenario trees that are inputted to multistage stochastic programming algorithms, as done in [100, 101]. Second, we can use GridFM–v0 and GridFMs in general as digital twins of power grids within Reinforcement Learning (RL) environments. Leveraging digital twins (data-driven representations of systems) enables faster computation of state transitions during RL training. This also allows RL agents to be trained on systems with highly complex dynamics that are difficult to model explicitly, such as distribution networks. As demonstrated in [102], a digital twin of a distribution network was used to train an RL agent, improving decision-making in both planning and operations. Third, GridFMs can also be used as a pre-trained representation for RL agents, for which FMs have been shown to be a valuable asset [103].
@@
The pre-training of GridFM–v0, which involves reconstructing node-level masked power flow data, can be complemented by edge masking and reconstruction. Extending this approach further, we will leverage the model’s generative capabilities to also propose new grid topologies that optimize for capacity expansion. By integrating constraints such as network reliability, load growth forecasts, and economic cost factors, GridFM can generate topologies that minimize congestion, maximize redundancy, and ensure efficient power distribution. Optimizing grid topology and resource dispatch co-optimization has already been achieved with GNNs in [104].
@@
GridFMs can be fine-tuned for anomaly detection tasks to identify patterns that signal potential cyber intrusions or physical failures [105]. We can also use GridFMs to identify vulnerabilities and conditions that lead to cascading failures [106].
Although the use cases mentioned above are the most apparent, experience with FMs and LLMs shows that once a model is available, numerous new applications can emerge. Therefore, the reader is encouraged to begin considering these possibilities, as FMs for the power grid will soon become a reality.
4.2.2 Building multi-modal GridFMs
Ideally, GridFMs should be able to handle more than one set of application tasks. This can be done by combining multiple GridFMs together [103] or by building self-contained stand-alone larger models, ingesting multi-modal spatial, temporal, as well as text data [14]. Finally, concepts such as Mixture-of-Experts (MoE) frameworks [107] can be exploited to compose GridFMs from a family of GridFMs.
Regardless of the method used, GridFMs will need to handle diverse data modalities in the long term, to account for complex external factors—such as climate, weather, economic and societal influences, distributed power sources, and the electrification of mobility and heating [76]. Training multi-modal FMs requires creating tokens for each modality, aligning them in space and time, and potentially normalizing or reducing their dimensionality [108]. Consequently, tasks beyond power flow and grid topology reconstruction [109] will need to be incorporated for pre-training. This approach has been demonstrated with models like 4M [110], which unifies different modalities into discrete tokens for computer vision. We anticipate that the number of trainable parameters for such models could approach those of weather models, reaching hundreds of millions [56, 111].
5 Conclusion and Discussion
@@
The rapid emergence of FMs is reshaping the AI landscape. While FMs are well established for text and image analysis, here we have introduced the idea of GridFMs: FMs for the power grid, able to learn from grid-relevant data to improve power grid operations, planning, and control. We first highlighted the strengths of FMs, which include excellent token prediction skills (signal reconstruction and forecasting), broad adaptability and scalability, and simulation acceleration. To demonstrate that GridFMs are well suited to close the growing gaps between computational capabilities and the industry’s need to cope with the increasing uncertainty and complexity of the power grid, we draw on the experiences of FMs for weather forecasting. The latter is a domain characterized by apparently intractable complexity. Nevertheless, we have seen scalable and adaptable FMs to produce much more computationally efficient models by more than 4 orders of magnitude than traditional weather forecasting models.
We have outlined concrete implementation phases to create an initial version, GridFM–v0, which is pre-trained on power flow data and tailored to accelerate power flow-related downstream applications (e.g., optimal power flow and contingency analysis). In the same way that FMs have been able to learn the fundamental equations governing weather, we envision GridFM–v0 to learn power flow, thereby creating a much more efficient model. The key advantage of such a GridFM–v0 is that once pre-trained on multiple grids and various networks and operating conditions, it can be readily fine-tuned to specific grid topologies and user requirements based on their own private data.
This paper advocates for research and development of FMs for the power grid, to facilitate the energy transition through collaboration between the various stakeholders. All stakeholders will benefit from GridFMs, simply by fine-tuning it to their specific needs and with their proprietary data and knowledge – without re-developing a full bespoke AI model each time. Towards that end, we envision, under the leadership of industry associations and open source initiatives, for the community of stakeholders to join forces and make the required investments to research and develop GridFMs in open source, thereby revolutionizing power grid analysis and leveraging AI advancements to meet common goals and needs.
@@
  • 1.
    H.F.H.: Conceptualization, Investigation, Writing – Original Draft, Project Administration
  • 2.
    B.G., L.S.A.M., A.P., A.V., J.W.: Conceptualization, Investigation, Writing – Original Draft
  • 3.
    J.B., A.B.M., T.B., S.C., I.F., B.H., R.J., K.K., V.M., F.M., M.D.M., O.R., H.S., L.X., E.S.Y., A.Z.: Investigation, Writing – Original Draft
  • 4.
    A.J.B., R.J.B., B.P.B., J.S., S.S.: Writing – Review & Editing
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Hendrik F. Hamann (hehamann.hh@gmail.com).
Acknowledgment
The work was supported by the respective organizations of the authors. The authors acknowledge the invaluable discussions and interactions in the working group for Foundation Models for the Power Grid, including Roya Ahmadi, Rich Argentieri, Gianni Vittorio Armani, Venkat Balachandran, Ranjini Bangalore, Venkat Banunarayanan, Edmon Begoli, Larry Bekkedahl, Keith Benes, Michelle Berti, Budhu Bhaduri, Roberto Bianchessi, Gilbert Bindewald, Louise Blais, David Bromberg, Matthias Bucher, Eren Çam, Supriya Chinthavali, Julien Chosson, Joseph Collum, Magnus Cormack, Peter De Bock, Fei Ding, Kathy Duviella, Darlain Edeme, Yamshid Farhat, Mark Fenton, Marcus Freitag, Alirez Ghassemian, Tassos Golnas, Gavin Goodland, Luca Grella, John Grosh, Benois Grossin, Andy Hackett, Hagen Haentsch, Zeryai Hagos, Grahame Hay, Frank Hellmann, Andy Henton, Raffael Hilber, Dan Hilton, Gabriela Hug, Alex Hurley, Sorana Ionescu, Kevin Johnston, Gareth Jones, Adrian Kelly, Bernd Klockl, Teddy Kott, Benjamin Kroposki, Teja Kuruganti, Larry Light, Dalton Lunga, Matt Magill, Adrian Maldonado, Antoine Maret, Pete Marsh, Russel Mason, James McClean, John Messer, Nikki Militello, Jordan Murkin, Xabi Muruaga, Indu Nambiar, Christian Nauck, Caleb Northrop, Gabby Nyirjesy , Marco Padula, Patrick Panciatici, Raul Alfaro Pelico, Mark Petri, Pierre Pinson, Abhishek Pontis, Josh Power, Alan Pressman, Feng Qiu, Mark Rafferty, Pippa Robertson, Nicoletta Rocca, Caroline Rose-Newport, Lyndon Ruff, Bryan Sacks, Kelly Sanders, Giovanni Sansavini, Michael Sheppard, Robert Shorten, Vikas Singhvi, Pratik Sonthalia, Brian Spears, Declan Stock, Olle Sundstroem, Matthew Tarduogno, Roger Thompson, Anna Carolina Tortora, Alex Vivas, Lucy Vu, Ben Weiner, Casey Werth, Daniel Wetzel, Gary White, Xin Wu, Steven Wysmuller, Yunchi Yang, Thomas Zadlo, Marzia Zafar and many others. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.